A sleigh carrying a load of presents weighs 9125 pounds. The same sleigh carrying a load of candy weights 7875 pounds. The toys weigh three times as much as the candy. How many pounds does the sleigh weigh when it is empty?
Answer by user_dude2008(1862)

s+p=9125
s+c=7875
p=3c

s+3c=9125
s+c=7875
2c=1250
c=625
s+c=7875
s+625=7875
s=7250

sleigh weighs 7250